OpenStudy (anonymous):

Given the word ALABAMA, how many distinguishable permutation or orders exist? A)3!/7! B)4!/7! C)7!/3! D)7!/4!

OpenStudy (anonymous):

if all the letters were different it would be \(7!\) right?

OpenStudy (anonymous):

D

OpenStudy (anonymous):

but since the word 'A' is used 4 times it would be 7!/4!

OpenStudy (anonymous):

but since there are 4 indistinguishable letters (A) it is \[\frac{7!}{4!}=7\times 6\times 5=210\]

OpenStudy (anonymous):

I thought nPr = n! / ( n -r )!

OpenStudy (anonymous):

this is not \(nPr\) though you are not permuting some number out of the total

OpenStudy (anonymous):

If not Permutation, so it makes sense !